OpenStudy (anonymous):

Can anyone help me put these numbers from least to greatest? (the 4.6 has a line above the 6) 3√4, 5.125, 4 3/5, 4.6, √9

OpenStudy (wendy.ivette11714):

\[3 \sqrt4 = \sqrt 4 = 2 = 3*2 = 6\] \[5.125 = 5.125\] \[4 \frac{ 3 }{ 5 } = \frac{ 12 }{ 5 }= 2.4\] \[4.6 = 4.6\] \[\sqrt9 = 3\] So the answer choices now are: 6, 5.125, 2.4, 4.6, 3

OpenStudy (wendy.ivette11714):

What do u mean the 6 has a line over it?

OpenStudy (wendy.ivette11714):

@karliroo18

OpenStudy (wendy.ivette11714):

The line on top of 6 means: 6.66666....

OpenStudy (wendy.ivette11714):

So the answer choices now are: 6, 5.125, 2.4, 4.6666..., 3

OpenStudy (wendy.ivette11714):

Now put them from least to greatest @karliroo18